OpenStudy (anonymous):

Help A triangle can be formed with side lengths 5, 7 and 10. true or false

OpenStudy (nikato):

true u can check by adding any of the 2 sides and the sum should always be greater than the 3rd

OpenStudy (mathstudent55):

This is how you find out. In a triangle, the the sum of the lengths of any two sides is greater than the length of the third side. Add the lengths of each pair of sides and compare to the length of the third side.

OpenStudy (anonymous):

THANKS

OpenStudy (mathstudent55):

These are the three pairs of 2 sides: 5, 7 5, 10 7, 10 Now add them. 5 + 7 = 12, which is greater than 10 (check) 5 + 10 = 15, which is greater than 7 (check) 7 + 10 = 17, which is greater than 7 (check) Since all sums check out, the statement is true.